OpenStudy (anonymous):

You buy 5 tickets to a concert for you and your friends. You pay $170, including a $20 service charge. How much does one ticket cost? A. 2003-05-05-00-00_files/i0120000.jpg B. 20x + 5 = 170 C. 5(20 + x) = 170 D. 5x + 20 = 170

OpenStudy (anonymous):

c

OpenStudy (anonymous):

I mean d. d is the correct answer

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):



OpenStudy (anonymous):

Anything else?

OpenStudy (anonymous):

A farmer builds a rectangular pen with a perimeter of 116 feet. If the length is 5 feet more than the width, what is the width? A. w + 5 = 116 B. 4w + 10 = 116 C. 2w + 5 = 116 D. 2w + 10 = 116

OpenStudy (anonymous):

d because there are two widths and 2 lengths. 5*2 is 10 and then subtract 10 from 116, divide by 2 then you get your width.

OpenStudy (anonymous):

Anything else?

OpenStudy (anonymous):

Which of the following is not a step used to solve multi-step equations? A. identify values of variables in the equation from the written word problem B. finish by getting all variables and constants on the same side of the equation C. plug in known variables from the word problem into the equation D. translate the written phrases into an algebraic equation